prove using mathematical induction that n2 is less than 2n for n>=5

Grade:12

1 Answers

Aditya Gupta
4 years ago
For n≥5  you have n^2>2n+1 (this should not be hard to see) so if n^2
2^(n+1)= 2⋅2^n>2n^2>n^2+2n+1=(n+1)^2.
If you still have doubts feel free to ask.
kindly approve :=)